OpenStudy (anonymous):

The ellipse with x-intercepts (5, 0) and (-5, 0); y-intercepts (0, 3) and (0, -3). How do i do this?

OpenStudy (anonymous):

They want me to find the equation of the ellipses.

OpenStudy (anonymous):



OpenStudy (anonymous):

So y= (3,-3) and x = (5,-5) ?

OpenStudy (anonymous):

you know what the center is right?

OpenStudy (anonymous):

Zero right?

OpenStudy (anonymous):

well actually it is \((0,0)\)

OpenStudy (anonymous):

the general equation is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] where the center is \((h,k)\) but since your center is \((0,0)\) you know it is \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] so now you job is only to find \(a^2\) and \(b^2\)

OpenStudy (anonymous):

\(a^2\) is easy to find because of the orientation, you know that \(a\) is the larger of the two numbers, and since the vertices are \((-5,0)\) and \((5,0)\) you know \(a=5\) and so \(a^2=25\)

OpenStudy (anonymous):

to find \(b^2\) compute \(5^2-3^2\)

OpenStudy (anonymous):

ok. So 16?